Let X denote the random variable distribution that gives 2,3,4,5,6,7,8,9 and 10 in a deck


of cards. Find the probability of X.

1
Expert's answer
2022-04-07T03:43:19-0400

We have four cards each of the (2-10) in a deck. So

Probability is

P(2)=P(3)=P(4)=P(5)=P(6)=P(7)=P(8)=P(9)=P(10)=4/52=1/13

Probability of one of this card is

P(2-10)=9x4/52=9/13